log(2x-1) + log(x+2) = 1
Answer by stanbon(57984)

log(2x-1) + log(x+2) = 1
---
log[(2x-1)(x+2)] = 1
---
(2x-1)(x+2) = 10^1
2x^2 + 3x -11 = 0
---
Quadratic Formula:
x = [-3 +- sqrt(9-4*2*-11)]/4
---
x = [-3 +- sqrt(97)]/4
====================================
Cheers,
Stan H.